ISSN 1201-561X

[figure: T; A; C]

THEORY AND APPLICATIONS OF CATEGORIES

Volume 25, 2011

A category of quantum categories Dimitri Chikhladze 1

A remark about the Connes fusion tensor product Andreas Thom 38

Remarks on punctual local connectedness Peter Johnstone 51

Comparative smootheology Andrew Stacey 64

Monoidal functor categories and graphic Fourier transforms Brian J. Day 118

Reflective-coreflective equivalence Erik Bdos, S. Kaliszewski, and John Quigg 142

An embedding theorem for adhesive categories Stephen Lack 180

Model-categories of coalgebras over operads Justin R. Smith 189

A small observation on co-categories Peter LeFanu Lumsdaine 247 Higher categorified algebras versus bounded homotopy algebras David Khudaverdyan, Ashis Mandal, and Norbert Poncin 250

\begin{matrix}c \\ ◯ \end{matrix} The Editors of Theory and Applications of Categories 2011. Permission to copy for private use

granted.

Symmetry and Cauchy completion of quantaloid-enriched categories Hans Heymans and Isar Stubbe 276

Towards a homotopy theory of higher dimensional transition systems Philippe Gaucher 294

Covariant presheaves and subalgebras Ulrich Höhle 342

On involutive monoidal categories J.M. Egger 368

The Faà di Bruno construction J.R.B. Cockett and R.A.G. Seely 394

Semidirect products and crossed modules in varieties of right Ω-loops Edward B. Inyangala 426

Yoneda theory for double categories Robert Paré 436

Flows: cocyclic and almost cocyclic Michael Barr, John F. Kennison, and R. Raphael 480

Countable meets in coherent spaces with applications to the cyclic spectrum Michael Barr, John F. Kennison, and R. Raphael 508

On reflective-coreflective equivalence and associated pairs Erik Bdos, S. Kaliszewski, and John Quigg 533

Differential restriction categories J.R.B. Cockett, G.S.H. Cruttwell, and J. D. Gallagher 537

Symbolic dynamics and the category of graphs Terrence Bisson and Aristide Tsemo 614

THEORY AND APPLICATIONS OF CATEGORIES (ISSN 1201-561X) will disseminate articles that significantly advance the study of categorical algebra or methods, or that make significant new contributions to mathematical science using categorical methods. The scope of the journal includes: all areas of pure category theory, including higher dimensional categories; applications of category theory to algebra, geometry and topology and other areas of mathematics; applications of category theory to computer science, physics and other mathematical sciences; contributions to scientific knowledge that make use of categorical methods.

Articles appearing in the journal have been carefully and critically refereed under the responsibility of members of the Editorial Board. Only papers judged to be both significant and excellent are accepted for publication.

Full text of the journal is freely available in .dvi, Postscript and PDF from the journal’s server at http://www.tac.mta.ca/tac/ and by ftp. It is archived electronically and in printed paper format.

Subscription information. Individual subscribers receive by e-mail abstracts of articles by e-mail as they are published. To subscribe, send e-mail to tac@mta.ca including a full name and postal address. For institutional subscription, send enquiries to the Managing Editor.

Information for authors. The typesetting language of the journal is TEX, and L^{A}TEX 2ε is strongly encouraged. Articles should be submitted by e-mail directly to a Transmitting Editor. Please obtain detailed information on submission format and style files from http://www.tac.mta.ca/tac/.

Managing editor. Robert Rosebrugh, Mount Allison University: rrosebrugh@mta.ca

TEXnical editor. Michael Barr, McGill University: barr@math.mcgill.ca

Assistant TEX editor. Gavin Seal, Ecole Polytechnique Fédérale de Lausanne:

gavin seal@fastmail.fm

Transmitting editors.

Clemens Berger, Université de Nice-Sophia Antipolis, cberger@math.unice.fr Richard Blute, Université d’ Ottawa: rblute@uottawa.ca Lawrence Breen, Université de Paris 13: breen@math.univ-paris13.fr Ronald Brown, University of North Wales: ronnie.profbrown(at)btinternet.com

Valeria de Paiva: valeria.depaiva@gmail.com

Ezra Getzler, Northwestern University: getzler(at)northwestern(dot)edu Kathryn Hess, Ecole Polytechnique Fédérale de Lausanne : kathryn.hess@epfl.ch Martin Hyland, University of Cambridge: M.Hyland@dpmms.cam.ac.uk P. T. Johnstone, University of Cambridge: ptj@dpmms.cam.ac.uk Anders Kock, University of Aarhus: kock@imf.au.dk Stephen Lack, Macquarie University: steve.lack@mq.edu.au F. William Lawvere, State University of New York at Buffalo: wlawvere@buffalo.edu Tom Leinster, University of Glasgow, Tom.Leinster@glasgow.ac.uk Jean-Louis Loday, Université de Strasbourg: loday@math.u-strasbg.fr Ieke Moerdijk, University of Utrecht: moerdijk@math.uu.nl Susan Niefield, Union College: niefiels@union.edu Robert Paré, Dalhousie University: pare@mathstat.dal.ca Jiri Rosicky, Masaryk University: rosicky@math.muni.cz Giuseppe Rosolini, Università di Genova: rosolini@disi.unige.it Alex Simpson, University of Edinburgh: Alex.Simpson@ed.ac.uk James Stasheff, University of North Carolina: jds@math.upenn.edu Ross Street, Macquarie University: street@math.mq.edu.au Walter Tholen, York University: tholen@mathstat.yorku.ca Myles Tierney, Rutgers University: tierney@math.rutgers.edu Robert F. C. Walters, University of Insubria: robert.walters@uninsubria.it R. J. Wood, Dalhousie University: rjwood@mathstat.dal.ca